OpenStudy (aaronandyson):

A metallic sphere is melted an then recasted into128 cones each of radius 4 cm and height 8 cm.Find the radius of the sphere.

OpenStudy (aaronandyson):

@hartnn

OpenStudy (aaronandyson):

@hartnn

OpenStudy (aaronandyson):

@pitamar

OpenStudy (anonymous):

Can you find the volume of a single cone?

OpenStudy (aaronandyson):

\[\frac{ 1 }{3? }*\pi*r^2h\]

OpenStudy (anonymous):

Yes

OpenStudy (aaronandyson):

So now?

OpenStudy (anonymous):

Now take this volume of a single cone and multiply by 128, because you have 128 cones. What do you get?

OpenStudy (aaronandyson):

Okay,Lemme calculate.

OpenStudy (anonymous):

How is it going?

OpenStudy (aaronandyson):

134.04 times 128?

OpenStudy (anonymous):

Do not approximate it just yet =) The answer may include \(\pi\) and can be a fraction. Let's just form it. If you want, I can show you what I mean.

OpenStudy (aaronandyson):

Sure.

OpenStudy (anonymous):

I made a mistake. let me fix it =|

OpenStudy (anonymous):

We know we a single cone has radius of 4cm and height of 8cm: $$\text{cone_volume} = \frac{\pi (4)^2 \cdot 8}{3} = \frac{\pi \big(2^2\big)^2 \cdot 8}{3} = \frac{\pi 2^4 \cdot 2^3}{3} = \frac{\pi 2^7}{3}$$ We have 128 cones in total so: $$\text{cones_volume} = 128 \cdot \text{cone_volume} = 128 \cdot \frac{\pi 2^7}{3} = \\ = 2^7 \cdot \frac{\pi 2^7}{3} = \frac{\pi 2^{7 + 7}}{3} = \frac{\pi 2^{14}}{3} $$Ok?

OpenStudy (aaronandyson):

I didnt get the lst step.

OpenStudy (anonymous):

The first step?

OpenStudy (anonymous):

Do you understand why \(\big(2^2\big)^2 = 2^4\)? What about \(2^4 \cdot 2^3 = 2^7\)?

OpenStudy (aaronandyson):

last step.

OpenStudy (anonymous):

128 is basically \(2^7\) so: $$ 128 \cdot \frac{\pi 2^7}{3} =\frac{\pi 2^7 \cdot 128}{3} = \frac{\pi 2^7 \cdot 2^7}{3} = \frac{\pi 2^{7+7}}{3} = 128 \cdot \frac{\pi 2^{14}}{3} $$Which is very close to what happend with the \(2^4\) and \(2^3\) above.

OpenStudy (anonymous):

We do this to find the volume of all the cones together. If we know the volume of a single cone and we know that we have 128 of these, then we just multiply the single-cone volume by 128 to get the total volume of all the cones. Is it clear so far?

OpenStudy (aaronandyson):

128 is used twice why?

OpenStudy (anonymous):

copy paste mistake. Let me rewrite this. Look above you'll see I've only used it once =p

OpenStudy (anonymous):

$$ 128 \cdot \frac{\pi 2^7}{3} =\frac{\pi 2^7 \cdot 128}{3} = \frac{\pi 2^7 \cdot 2^7}{3} = \frac{\pi 2^{7+7}}{3} = \frac{\pi 2^{14}}{3} $$

OpenStudy (anonymous):

Here =)

OpenStudy (aaronandyson):

so 128^2*pi/3?

OpenStudy (anonymous):

yes, but I've written this as a power of 2 for a reason. So I prefer to stick with \(\frac{\pi 2^{14}}{3}\) for now, ok?

OpenStudy (aaronandyson):

so 16384/3 * 22/7?

OpenStudy (anonymous):

Well, again you try to approximate. Remember we are not asked about a volume or anything, we are asked about the radius of the sphere, which brings to the question "Why did we calculate the this volume then?". Well, if we assume that the sphere is now hollow (because it isn't said to be in the question) and that we didn't somehow change the density of the material it is made of and made it take less volume.... then we can say that the melting process didn't change the total volume of what we have. That means that the volume of our cones is the volume of our sphere. Do you know the volume formula for sphere?

OpenStudy (aaronandyson):

\[=\Pi \times r^2 \times h\]

OpenStudy (anonymous):

Well, sphere doesn't have any height so that cannot be true. Sphere only has radius. the formula is: $$ \text{sphere_volume} = \frac{4\pi r^3}{3} $$This has a neat calculus proof, if you know some calculus then maybe I can show you later =) Anyway, we know that our sphere volume is equal to our total cones volume. So: $$ \text{sphere_volume} = \text{cones_volume} \\ \frac{4\pi r^3}{3} = \frac{\pi 2^{14}}{3} $$Ok?

OpenStudy (aaronandyson):

n0w?

OpenStudy (aaronandyson):

r^2 = 2^14?

OpenStudy (anonymous):

Not exactly. We can multiply both sides by 3 to get rid of it for example. What else can we get rid of on both sides?

OpenStudy (aaronandyson):

pi

OpenStudy (anonymous):

Right. and now we're left with: $$ \frac{4\cancel \pi r^3}{\cancel3} = \frac{\cancel\pi 2^{14}}{\cancel3} \\ 4 r^3 = 2^{14} $$ We still have to get rid of the 4 on the left. So what is \(\frac{2^{14}}{4}\)?

OpenStudy (aaronandyson):

16384/4?

OpenStudy (anonymous):

Don't work so hard =) we can use power rules again in our favour So we have say \(4 = 2^2\) which makes it \(\frac{2^{14}}{2^2}\) What happens now?

OpenStudy (aaronandyson):

2^12?

OpenStudy (anonymous):

Exactly. So $$ r^3 = 2^{12} $$ Now what?

OpenStudy (aaronandyson):

4096?

OpenStudy (aaronandyson):

so 64^2 is 4096

OpenStudy (anonymous):

We have no need to change the form just yet. We are still in our way to find r and this form might actually help us. What is the next operation we do in order to find r?

OpenStudy (anonymous):

$$r^3 = 2^{12}$$

OpenStudy (anonymous):

Well, you don't have to compute it, just name it =) What is the last thing that we have to get rid of in order to find \(r\)?

OpenStudy (aaronandyson):

pi

OpenStudy (aaronandyson):

sorry sorry 4

OpenStudy (anonymous):

We don't have pi anymore (or 4), we have $$r^3 = 2^{12}$$ It's good thing that we know \(r^3\), but we want \(r\). What should we do to change the \(r^3\) to \(r\)?

OpenStudy (anonymous):

Hint?

OpenStudy (aaronandyson):

sqrt2^12

OpenStudy (anonymous):

Well we're getting closer. square root (\(\sqrt{\text{something}}\)) will get rid of a power of 2, but we have a power of 3, so we need a cube root: $$ r^3 = 2^{12} \\ \sqrt[3]{r^3} = \sqrt[3]{2^{12}} \\ r = \sqrt[3]{2^{12}} $$

OpenStudy (anonymous):

Now, if you remember power rules (again), then cube root is just like power of 1/3. So we can say: $$ \sqrt[3]{2^{12}} = \big(2^{12}\big)^{\frac{1}{3}} $$What happens now?

OpenStudy (aaronandyson):

2^4

OpenStudy (anonymous):

exactly. So r = 2^4 = ?

OpenStudy (aaronandyson):

64 which i had said 15 min back.

OpenStudy (aaronandyson):

Thanks anyways.

OpenStudy (anonymous):

\(2^4 = 64\)?

OpenStudy (aaronandyson):

16 sorry

OpenStudy (anonymous):

Right =) You're welcome